OpenStudy (zmudz):

Let \(f:\mathbb R \to \mathbb R\) be a function such that for any irrational number \(r\), and any real number \(x\) we have \(f(x)=f(x+r)\). Show that \(f\) is a constant function.

OpenStudy (zarkon):

what have you tried

OpenStudy (zarkon):

the proof is quite short

OpenStudy (zmudz):

I am having trouble even getting started. I don't know what I'm supposed to do. Any help is greatly appreciated.

OpenStudy (zarkon):

choose x=0 then f(0)=f(0+r)=f(r) for all r irrational so f is constant on the irrational numbers now you need to extend this to the rational numbers consider x=a-r where a is a rational number